# Flow Straighter and Faster: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories

### v1

@@ 4.2 Re-Meanflow: Efficient One-Step Generative Modeling via MeanFlow on Rectified Trajectories
Motivated by these observations, we propose Re-Meanflow, which models the mean velocity field along rectified trajectories, avoiding relying on perfectly straight ODE paths. Concretely, given a pretrained flow model $v_{\theta}$ trained on the independent couplings $p_{\mathbf{x}\mathbf{z}}=p_{\mathbf{x}}p_{\mathbf{z}}$, we obtain a straighter coupling distribution $p^{1}_{\mathbf{x}\mathbf{z}}$ by transporting samples using the 1-rectified flow $v_{\theta}^{1}$. We then train a MeanFlow model $u_{\theta}$ on these rectified couplings under the MeanFlow objective (Eq. 5).
This combination substantially accelerates MeanFlow training compared to using independent couplings. Rectified trajectories contain far fewer path intersections, allowing conditional velocities to better approximate the marginal velocity field and also keeping instantaneous velocities closer to their temporal averages. As a result, the discrepancy between $u(\mathbf{z}_{t},r,t)$ and $v(\mathbf{z}_{t},t)$ in Eq. 5 is significantly reduced.
Toy Evidence (Panel f). Re-Meanflow (ours) achieves more accurate generation and yields many fewer invalid one-step samplings, demonstrating the complementary roles of the two components.
Takeaway. Trajectory rectification sufficiently reduces curvature to enable efficient MeanFlow training, while MeanFlow removes the need for fully straight trajectories. This synergistic advantage of Re-Meanflow is further evident on real ImageNet data in Fig. 1, where it outperforms using MeanFlow or Rectified Flow alone. A visual illustration of this synergy is provided in Fig.2.
4.3 Distance Truncation
Figure 4: Distance-Error Correlation. Histogram of data-noise $\ell_{2}$ distances on ImageNet $512^{2}$, colored by angular error (computed as $\arccos\big(\tfrac{\langle\mathbf{z}-\mathbf{x},\;u_{\theta}(\mathbf{z},0,1)\rangle}{\|\mathbf{z}-\mathbf{x}\|_{2}\,\|u_{\theta}(\mathbf{z},0,1)\|_{2}}\big)$, the angle between the predicted and ground-truth velocities). Errors are computed using a MeanFlow model trained on untruncated 1-rectified couplings (config (d) in Tab. 3). A clear high-distance high-error tail (90th percentile marked) motivates our Distance Truncation.
As discussed in Sec. 4.1, a small but consequential subset of couplings exhibits large data-noise distances, often arising from low-density regions where samples must travel substantially farther than typical cases. Such pairs tend to produce higher curvature even after rectification, contributing disproportionately to instability during training.
This observation motivates a simple filtering heuristic: after generating candidate couplings, we discard the top $k\%$ (e.g., $10\%$) with the largest data-noise $\ell_{2}$ distance. This Distance Truncation selectively removes the pairs most associated with excessive transport length, and in practice leads to noticeably straighter flows and more stable learning.
We also examine this behavior on real datasets. As shown in Fig. 4, couplings with large data-noise distances consistently populate the high-loss tail of the flow regression objective, suggesting that they coincide with high-curvature, high-variance regions. As shown in our empirical results in Sec. 5.2, filtering these pairs yields a significant performance gain in practice.
@@ 5 Experiments
Datasets. We conduct experiments on ImageNet [7] at $64^{2}$ resolution in pixel space and at $256^{2}$ and $512^{2}$ resolutions in latent space [36].
Settings. We initialize Re-Meanflow from a pretrained flow or diffusion model. For ImageNet-$64^{2}$ and ImageNet-$512^{2}$, we initialize Re-Meanflow from the pretrained EDM2-S model [23]. For ImageNet-$256^{2}$, we initialize from the pretrained SiT-XL model [31]. During the reflow process, we generate 5M data-noise couplings using the default procedures described in the corresponding papers. During sampling for better synthetic image quality, we apply classifier-free guidance (CFG)[16] for ImageNet-$256^{2}$ and Autoguidance[22] for ImageNet-$64^{2}$ and $512^{2}$. More details on the implementation and the hyperparameter settings are provided in Appendix B.
Baselines. We compare Re-Meanflow against recent state-of-the-art one-step flow-based methods, selecting baselines with comparable architecture or computational cost to ensure fair comparisons.
5.1 One Step Generation Quality
For all experiments, we evaluate image quality using Fréchet Inception Distance (FID) [15]. When comparing with prior methods, we select those with comparable architectures and parameter counts. Across all evaluated resolutions, Re-Meanflow consistently outperforms state-of-the-art one-step flow-based generation methods.
On the EDM2-S backbone [23], Re-Meanflow achieves superior performance at both $64^{2}$ and $512^{2}$ resolutions. On ImageNet-$64^{2}$, Re-Meanflow outperforms the closely related 2-rectified flow++ [24] by 33.4% in FID, and slightly improves over recent state-of-the-art one-step baselines[29, 25, 37]. On ImageNet-$512^{2}$, Re-Meanflow also delivers strong one-step image quality, achieving a 9% FID gain over AYF[37] and outperforming strong consistency distillation methods[37, 19].
On the SiT backbone [31] for ImageNet-$256^{2}$, Re-Meanflow slightly surpasses MeanFlow [11] despite being trained without real-image supervision and relying solely on synthesized reflow samples. This is noteworthy because training exclusively on self-generated data is known to degrade performance due to self-conditioning effects [2]. We attribute this improvement to the more favorable optimization landscape created by rectified mean-velocity learning: although the supervision is limited to synthetic couplings, rectification produces a smoother and lower-variance trajectory family, effectively narrowing the search space. As a result, the model converges more reliably toward a high-quality solution.
Table 1: Class-conditional generation on ImageNet. All results use classifier-free guidance (CFG) when supported; $\times 2$” indicates that CFG doubles the effective NFE. denotes methods initialized from, or directly comparable to, the diffusion backbones also marked with . (iCT[42] result at $256^{2}$ is reported by IMM[57], and ECT/ECD[10] results at $512^{2}$ are reported by from CMT[19]).
@@ 5.1 One Step Generation Quality
EDM2-S [23] $63\times 2$ 1.58
   + Autoguidance [22] $63\times 2$ 1.01
EDM2-XL [23] $63\times 2$ 1.33
@@
2-rectified flow++ [24] 1 4.31
iCT [42] 1 4.02
ECD-S [10] 1 3.30
sCD-S [29] 1 2.97
TCM [25] 1 2.88
AYF [37] 1 2.98
@@
ADM [8] $250\times 2$ 10.94
DiT-XL [34] $50\times 2 2.2$7
SiT-XL [31] $250\times 2$ 2.06
SiT-XL + REPA  [51] $250\times 2$ 1.42
@@
iCT [42] 1 34.6
SM [9] 1 10.6
iSM [32] 1 5.27
iMM [57] $1\times 2$ 7.77
MeanFlow [11] 1 3.43
@@
EDM2-S [23] $63\times 2$ 2.23
   + Autoguidance [22] $63\times 2$ 1.34
EDM2-XXL [23] $63\times 2$ 1.81
@@
ECT [10] 1 9.98
ECD [10] 1 8.47
CMT [19] 1 3.38
sCT-S [29] 1 10.13
sCD-S [29] 1 3.07
AYF [37] 1 3.32
@@
Figure 5: Qualitative results for Re-Meanflow (NFE=1) on ImageNet $64^{2}$ (Left), $256^{2}$ (Middle), and $512^{2}$ (Right).
@@ 5.2 Training Efficiency
As noted by Lee et al. [24], distillation methods involving reflow can remain computationally competitive, even though they require generating additional data-noise couplings.
Synergistic Benefits of Re-Meanflow.
Figure 6: Comparison of convergence on ImageNet-$256^{2}$. Re-Meanflow achieves the best FID under the same training budget.
We compare Re-Meanflow against MeanFlow [11] and 2-rectified flow++ [24] on ImageNet-$512^{2}$ without using CFG. All methods are initialized from EDM2-S [23] and follow the recommended training configurations. Re-Meanflow and 2-rectified flow++ are trained on the same set of 5M couplings generated during the reflow process, whereas MeanFlow is trained on independent couplings.
For fairness, we follow the training settings reported in the original papers with one exception: we do not include the LPIPS [54] loss for 2-rectified flow++. LPIPS operates in pixel space and cannot be used directly in the latent space setting of our experiments; moreover, LPIPS may leak perceptual information correlated with FID, potentially exaggerating improvements. Instead, we adopt the Huber loss, the strongest configuration reported for 2-rectified flow++ without LPIPS.
@@
This demonstrates that combining MeanFlow with rectified flow is not merely an additive combination of two components: rectification reduces trajectory variance and stabilizes the MeanFlow objective, while mean-velocity learning enables efficient one-step sampling. The two components reinforce each other, yielding a synergistic effect that neither method achieves independently.
Overall Training Cost Comparison.
@@
We further evaluate Re-Meanflow by estimating the total training cost in FLOPs and GPU hours on ImageNet-$64^{2}$. The protocol for computing these metrics follows Lee et al. [24] and is detailed in the Appendix C.1. We compare against AYF [37], the strongest existing distillation baseline in both quality and efficiency, and against the closely related 2-rectified flow++ [24].

### v2

@@ 3.2 Taming the Curvature Bottleneck
To construct this coupling that induces straighter trajectories, we follow Rectified Flow [liu2022flow] and perform a single reflow step using a pretrained flow model $v_{\phi}$ trained under the independent coupling. Concretely, we sample $\mathbf{z}\sim p_{\mathbf{z}}$ and define $\mathbf{x}\;=\;\mathbf{z}-\int_{0}^{1}v_{\phi}(\mathbf{z}_{\tau},\tau)\,d\tau,$ which induces a new coupling distribution $p^{1}_{\mathbf{x}\mathbf{z}}(\mathbf{x},\mathbf{z})$. We then train Re-MeanFlow by modeling the mean-velocity field on $p^{1}_{\mathbf{x}\mathbf{z}}$.
3.3 Estimating the Curvature of the Generative Trajectory
To quantify the trajectory curvature over time pairs $(r,t)$ , we propose an angle-based proxy. Let $\tilde{u}(\mathbf{z}_{t},r,t)$ denote the mean-velocity over the interval $[r,t]$. In practice, we approximate this quantity by integrating a learned velocity field $v_{\theta}$ trained under the corresponding coupling:
$\tilde{u}(\mathbf{z}_{t},r,t)\;=\;\frac{1}{t-r}\int_{r}^{t}v_{\theta}(\mathbf{z}_{\tau},\tau)\,d\tau,$ (6)
and let $v_{\theta}(\mathbf{z}_{t},t)$ denote the instantaneous velocity at $(\mathbf{z}_{t},t)$. We then define
$\mathrm{Curv}(r,t)\;=\;\mathbb{E}_{\mathbf{z}_{t}\sim p_{t}}\Big[\angle\!\big(\tilde{u}(\mathbf{z}_{t},r,t),\,v_{\theta}(\mathbf{z}_{t},t)\big)\Big],$ (7)
where $\angle(\mathbf{a},\mathbf{b})=\arccos\!\left(\frac{\langle\mathbf{a},\mathbf{b}\rangle}{\|\mathbf{a}\|\,\|\mathbf{b}\|}\right)$. Intuitively, $\mathrm{Curv}(r,t)$ measures directional disagreement between (i) the average transport direction over the interval $[r,t]$, given by $\tilde{u}(\mathbf{z}_{t},r,t)$, and (ii) the instantaneous flow direction at time $t$, given by $v_{\theta}(\mathbf{z}_{t},t)$. This proxy serves as a practical diagnostic similar to the $\ell_{2}$-based curvature measures used in Rectified Flow [liu2022flow]; in particular, for perfectly straight trajectories, we have $\mathrm{Curv}(r,t)=0$ for all valid $(r,t)$. We visualize $\mathrm{Curv}(r,t)$ in the ImageNet-$512^{2}$ over the $(r,t)$-plane in Fig. 2 and show that the rectified couplings used by Re-MeanFlow exhibit substantially lower trajectory curvature. More results at other resolutions are provided in Appendix 0.C.4.
Algorithm 1 Re-MeanFlow
1:Input: pretrained rectified flow $v_{\phi}$; prior $p_{\mathbf{z}}$; ODE solver SamplePair; time distribution $p_{r,t}$; guidance strength $\omega$; #pairs NN; truncation ratio $k\%$; training iterations $T$; learning rate $\eta$.
2:Output: trained Re-MeanFlow model $u_{\theta}$.
3:Initialize $\theta\leftarrow\textsc{InitFrom}(\phi)$ $\triangleright$ Initialize $u_{\theta}$ from teacher $v_{\phi}$
4:Stage A: Sample rectified couplings
[…]
6:Compute distances $d_{i}\leftarrow\|\mathbf{x}_{i}-\mathbf{z}_{i}\|_{2}$ for all pairs
[…]
9:Stage B: Train Re-MeanFlow on $\mathcal{P}_{\text{keep}}$
10:for $s=1,\dots,T$ do
11:  Sample $(\mathbf{x},\mathbf{z})\sim\mathcal{P}_{\text{keep}}$, $(r,t)\sim p_{r,t}$
12:  $\mathbf{z}_{t}\leftarrow(1-t)\mathbf{x}+t\mathbf{z}$
13:  $u_{\text{tgt}}\leftarrow(\mathbf{z}-\mathbf{x})-(t-r)\,\frac{d}{dt}u_{\theta}(\mathbf{z}_{t},r,t)$
14:  $\mathcal{L}_{\text{MF}}(\theta)\leftarrow\big\|u_{\theta}(\mathbf{z}_{t},r,t)-\operatorname{sg}(u_{\text{tgt}})\big\|_{2}^{2}$
15:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{MF}}(\theta)$
16:end for
17:Stage C: CFG fine-tuning
18:Fine-tune $u_{\theta}$ on $v^{\text{cfg}}_{\omega}$ (Eq. 8) using the loop in Stage B.
19:Return $u_{\theta}$
3.4 Training
Following prior work [liu2022flow, lee2024improving], we first construct the rectified coupling distribution $p^{1}_{\mathbf{x}\mathbf{z}}(\mathbf{x},\mathbf{z})$ by generating $N$ data-noise pairs with a pretrained flow model $v_{\phi}$, to ensure sufficient coverage of both the data and noise distributions.
3.4.1 Classifier-free guidance.
To enable classifier-free guidance (CFG) at inference, Re-MeanFlow (as in MeanFlow) must be trained to predict the CFG mean-velocity field $v_{\omega}^{\text{cfg}}$:
$v^{\text{cfg}}_{\omega}(\mathbf{z}_{t},t\mid c)\triangleq\omega\,v(\mathbf{z}_{t},t\mid c)+(1-\omega)\,v(\mathbf{z}_{t},t),$ (8)
where $\omega$ denotes the guidance strength and $c$ is the conditioning signal. In practice, we find it more stable to apply guidance via a two-stage procedure: we first train without guidance to initialize the model, and then perform a brief fine-tuning stage using the CFG objective. Overall, Re-MeanFlow converges rapidly, typically within a single pass over the generated couplings.
Figure 3: Left: Distance statistics of the data-noise pair distance $\|\mathbf{x}-\mathbf{z}\|_{2}$ under the independent coupling $p^{0}_{\mathbf{x}\mathbf{z}}$ (red) and the rectified coupling $p^{1}_{\mathbf{x}\mathbf{z}}$ (green). Right: Histogram of $\|\mathbf{x}-\mathbf{z}\|_{2}$ for rectified pairs $(\mathbf{x},\mathbf{z})\!\sim\!p^{1}_{\mathbf{x}\mathbf{z}}$, with bins colored by the curvature proxy (Sec. 3.3). Overall, rectification reduces the average pair distance between data and noise, but a small long-distance tail remains; these long-distance couplings tend to exhibit higher curvature, motivating our distance-based truncation to remove many high-curvature pairs and further improve training stability and performance.
3.4.2 Distance-based truncation.
Rectified Flow [liu2022flow] guarantees that rectification does not increase transport cost: for any convex cost $c:\mathbb{R}^{d}\!\to\!\mathbb{R}$,
$\mathbb{E}_{p^{1}_{\mathbf{x}\mathbf{z}}}\!\left[c(\mathbf{x},\mathbf{z})\right]\;\leq\;\mathbb{E}_{p^{0}_{\mathbf{x}\mathbf{z}}}\!\left[c(\mathbf{x},\mathbf{z})\right].$ (9)
We empirically verify this with $c(\mathbf{x},\mathbf{z})=\|\mathbf{x}-\mathbf{z}\|_{2}$ and plot the resulting distance distributions in Fig. 3 (left). As expected, the rectified coupling $p^{1}_{\mathbf{x}\mathbf{z}}$ is left-shifted relative to the original coupling $p^{0}_{\mathbf{x}\mathbf{z}}$, indicating smaller transport distances. Notably, however, $p^{1}_{\mathbf{x}\mathbf{z}}$ still exhibits a non-negligible long-tail subset of rectified pairs $(\mathbf{x},\mathbf{z})\!\sim\!p^{1}_{\mathbf{x}\mathbf{z}}$ have unusually large $\|\mathbf{x}-\mathbf{z}\|_{2}$ (above the dashed 90th-percentile threshold). We find that these long-distance pairs are also systematically more curved under our curvature proxy (Sec. 3.3), as shown in Fig. 3 (right). Motivated by this correlation, we introduce a simple distance-based truncation heuristic: we discard rectified couplings whose $\|\mathbf{x}-\mathbf{z}\|_{2}$ exceeds a fixed percentile threshold. In practice, we found that truncating the top $\mathbf{10\%}$ largest-distance couplings works robustly across all settings.
3.4.3 Algorithm.
The full algorithm of Re-MeanFlow is provided in Alg.1.
@@ 4 Experiments
4.0.1 Datasets.
We conduct experiments on ImageNet [deng2009imagenet] at $64^{2}$ resolution in pixel space and at $256^{2}$ and $512^{2}$ resolutions in latent space [rombach2022high].
4.0.2 Settings.
We initialize Re-MeanFlow from a pretrained flow or diffusion model. For ImageNet-$64^{2}$ and ImageNet-$512^{2}$, we initialize Re-MeanFlow from the pretrained EDM2-S model [karras2024analyzing]. For ImageNet-$256^{2}$, we initialize from the pretrained SiT-XL model [ma2024sit]. To rectify the trajectory, we generate 5M data-noise couplings using the default procedures described in the corresponding papers. During sampling for better synthetic image quality, we apply classifier-free guidance (CFG)[ho2022classifier] for ImageNet-$256^{2}$ and Autoguidance[karras2024guiding] for ImageNet-$64^{2}$ and $512^{2}$. More details on the implementation and the hyperparameter settings are provided in Appendix  0.B.
4.0.3 Baselines.
We compare Re-MeanFlow against recent state-of-the-art one-step flow-based methods, selecting baselines with comparable architecture or computational cost to ensure fair comparisons.
Table 1: Class-conditional generation on ImageNet. All results use classifier-free guidance (CFG) when supported; $\times 2$” indicates that CFG doubles the effective NFE. denotes methods initialized from, or directly comparable to, the diffusion backbones also marked with . (iCT[song2023improved] result at $256^{2}$ is reported by IMM[zhou2025inductive], and ECT/ECD[geng2024consistency] results at $512^{2}$ are reported by CMT[hu2025cmt]). Re-MeanFlow (ours) achieves the best FID across all settings, outperforming prior state-of-the-art one-step distillation and training-from-scratch methods.
@@
EDM2-S [karras2024analyzing] $63\times 2$ 1.58
+ Autoguidance [karras2024guiding] $63\times 2$ 1.01
EDM2-XL [karras2024analyzing] $63\times 2$ 1.33
@@
2-rectified flow++ [lee2024improving] 1 4.31
iCT [song2023improved] 1 4.02
ECD-S [geng2024consistency] 1 3.30
sCD-S [lu2024simplifying] 1 2.97
TCM [lee2024truncated] 1 2.88
AYF [sabour2025align] 1 2.98
@@
ADM [dhariwal2021diffusion] $250\times 2$ 10.94
DiT-XL [peebles2023scalable] $50\times 2 2.2$7
SiT-XL [ma2024sit] $250\times 2$ 2.06
SiT-XL + REPA [yu2024representation] $250\times 2$ 1.42
@@
iCT [song2023improved] 1 34.6
SM [frans2024one] 1 10.6
iSM [nguyen2025improved] 1 5.27
iMM [zhou2025inductive] $1\times 2$ 7.77
MeanFlow [geng2025mean] 1 3.43
@@
EDM2-S [karras2024analyzing] $63\times 2$ 2.23
+ Autoguidance [karras2024guiding] $63\times 2$ 1.34
EDM2-XXL [karras2024analyzing] $63\times 2$ 1.81
@@
ECT [geng2024consistency] 1 9.98
ECD [geng2024consistency] 1 8.47
CMT [hu2025cmt] 1 3.38
sCT-S [lu2024simplifying] 1 10.13
sCD-S [lu2024simplifying] 1 3.07
AYF [sabour2025align] 1 3.32
@@
Figure 4: Visualization of Re-MeanFlow one-step generation (NFE=1) on ImageNet at $64^{2}$ (Left), $256^{2}$ (Middle), and $512^{2}$ (Right).
4.1 One-Step Generation Quality
For all experiments, we evaluate image quality using Fréchet Inception Distance (FID) [heusel2017gans] in Tab.1. When comparing with prior methods, we select those with comparable architectures and parameter counts. Across all evaluated resolutions, Re-MeanFlow consistently outperforms state-of-the-art one-step flow-based generation methods. The qualitative results are shown in Fig. 4
On the EDM2-S backbone [karras2024analyzing], Re-MeanFlow achieves superior performance at both $64^{2}$ and $512^{2}$ resolutions. On ImageNet-$64^{2}$, Re-MeanFlow outperforms the closely related 2-rectified flow++ [lee2024improving] by 33.4% in FID, and slightly improves over recent state-of-the-art one-step baselines[lu2024simplifying, lee2024truncated, sabour2025align]. On ImageNet-$512^{2}$, Re-MeanFlow also delivers strong one-step image quality, achieving a 9% FID gain over AYF[sabour2025align] and outperforming strong consistency distillation methods[sabour2025align, hu2025cmt].
On the SiT backbone [ma2024sit] for ImageNet-$256^{2}$, Re-MeanFlow slightly surpasses MeanFlow [geng2025mean] despite being trained without real-image supervision and relying solely on synthesized samples. This is noteworthy because training exclusively on self-generated data is known to degrade performance due to self-conditioning effects [alemohammad2023self]. We attribute this improvement to the more favorable optimization landscape created by rectified mean-velocity learning: although the supervision is limited to synthetic couplings, rectification produces a smoother and lower-variance trajectory family. As a result, the model converges more reliably toward a high-quality solution.
Figure 5: Re-MeanFlow yields faithful one-step sampling with significantly less training. (a) Multi-step samples $\mathbf{x}_{\text{mul}}$ (reference) obtained by integrating the predicted mean-velocity field; multi-step results from MeanFlow and Re-MeanFlow are visually similar, and we show a mixed set from both methods. (b) MeanFlow one-step samples $\mathbf{x}_{\text{one}}$ after 20k iterations remain blurry and deviate noticeably from $\mathbf{x}_{\text{mul}}$. (c) Re-MeanFlow one-step samples after 10k iterations are already sharp and closely match $\mathbf{x}_{\text{mul}}$. (d) Sampling discrepancy distribution $\|\mathbf{x}_{\text{mul}}-\mathbf{x}_{\text{one}}\|$: Re-MeanFlow is strongly left-shifted, indicating substantially smaller one-step vs. multi-step mismatch. All results are on ImageNet $512^{2}$, using the same noise seeds for both methods.
@@ 4.2 Training Efficiency
4.2.1 Compared to MeanFlow.
We compare Re-MeanFlow against MeanFlow on ImageNet $512^{2}$ without classifier-free guidance (CFG). As shown in Fig. 1, Re-MeanFlow converges substantially faster: even when MeanFlow is trained with $2\times$ the compute budget, its one-step samples remain noticeably blurry, whereas Re-MeanFlow already produces sharp images, with a large FID gap (8.6 vs. 30.9). To diagnose this difference, we contrast one-step samples $\mathbf{x}_{\mathrm{one}}=\mathbf{z}-u_{\theta}(\mathbf{z},0,1)$ against multi-step samples $\mathbf{x}_{\mathrm{mul}}=\mathbf{z}-\int_{0}^{1}u_{\theta}(\mathbf{z},\tau,\tau)d\tau$ obtained by simulating the integral of the instantaneous-velocity (Fig. 5). Both methods are initialized from the same EDM2-S checkpoint, while MeanFlow is trained with $2\times$ the compute budget. Qualitatively, Re-MeanFlow produces nearly indistinguishable samples under one-step and multi-step generation, indicating that $u_{\theta}(\mathbf{z},0,1)$ closely matches the multi-step estimate. In contrast, MeanFlow’s one-step predictions are significantly degraded. Quantitatively, we measure the sampling discrepancy as $\|\mathbf{x}_{\mathrm{mul}}-\mathbf{x}_{\mathrm{one}}\|$ and plot its distribution in Fig. 5d. Re-MeanFlow exhibits a pronounced left shift, confirming a smaller one-step-to-multi-step error and hence more accurate mean-velocity modeling of the underlying trajectories. We provide additional convergence comparisons with MeanFlow and other related methods in Appendix 0.C.3.
4.2.2 Compared to Other One-Step Distillation Methods.
As noted by [lee2024improving], distillation methods involving sampling rectified couplings can still remain computationally competitive in overall computation cost, even though they require generating additional data-noise couplings.
@@
We evaluate Re-MeanFlow by estimating the total computation cost in FLOPs and GPU hours on ImageNet-$64^{2}$. The protocol for computing these metrics follows [lee2024improving] and is detailed in the Appendix 0.C.2. We compare against AYF [sabour2025align], the strongest existing distillation baseline in both quality and efficiency, and against the closely related 2-rectified flow++ [lee2024improving].
@@
We also compare estimated FLOPs. Although Re-MeanFlow shows a smaller advantage under this metric, FLOPs alone do not fully reflect practical runtime. A substantial portion of our compute lies in the inference-only coupling sampling stage, which can be executed on widely accessible inference-grade GPUs and runs significantly faster than training workloads with the same FLOP count.
Figure 7: Loss landscape of $\mathcal{L}_{MF}$ on a PCA plane of $\mathbf{z}_{t}$. We visualize $\mathcal{L}_{MF}$ as a function of the input $\mathbf{z}_{t}$ by evaluating the loss on a 2D grid spanned by the top two PCA directions of a batch of $\mathbf{z}_{t}$ samples. We fix $t=0.7$ and sweep $r\in\{0.7,0.5,0.3,0.1\}$ (left to right), where larger trt-r corresponds to a harder, longer-jump $(r,t)$ regression task. Top (MeanFlow): the landscape becomes increasingly sharp and irregular as trt-r grows. Bottom (Re-MeanFlow; ours): the landscape remains substantially smoother and better-conditioned across all $r$. Overall, learning mean velocity on straighter trajectories yields a markedly smoother objective, enabling more efficient and stable training.
4.3 Loss Landscape
To explain the training speedup enabled by learning mean-velocity on straighter trajectories, we visualize in Fig. 7 the loss landscape of $\mathcal{L}_{MF}$ on ImageNet $256^{2}$ with respect to the input $\mathbf{z}_{t}$ for both MeanFlow and Re-MeanFlow. We fix $t=0.7$ and plot the landscape for different values of $r$. Additional results at other resolutions are provided in Appendix 0.C.1.
To construct these visualizations, we perform PCA on a batch of $\mathbf{z}_{t}$ samples to obtain the top two principal directions. We then evaluate $\mathcal{L}_{MF}$ on a 2D grid spanned by these directions and centered at the sampled $\mathbf{z}_{t}$, and visualize the resulting surfaces. As shown in Fig. 7, Re-MeanFlow consistently exhibits a smoother and better-conditioned landscape. In particular, as trt-r grows, the MeanFlow objective becomes increasingly sharp and irregular, whereas Re-MeanFlow remains substantially smoother and more stable.
Unlike conventional loss-landscape studies [li2018visualizing] that perturb model parameters, we perturb $\mathbf{z}_{t}$ to directly measure the conditioning of the mean-velocity regression induced by the trajectory geometry. When trajectories are highly curved, the interval mean velocity becomes a rapidly varying, locally unstable function of $\mathbf{z}_{t}$, yielding a rugged loss landscape in $\mathbf{z}_{t}$-space. Such ruggedness indicates that the regression target itself is poorly conditioned, which in turn hampers optimization. In contrast, rectified (straighter) trajectories markedly smooth the landscape as a function of $\mathbf{z}_{t}$, producing a better-conditioned learning problem and enabling more efficient training.
Table 2: Effect of distance-based truncation strength (discarding a top fraction of couplings by $\ell_{2}$ endpoint distance). We report FID/IS and Precision/Recall on ImageNet $512^{2}$ (top) and $256^{2}$ (bottom). Mild truncation consistently improves stability and sample quality (lower FID, higher IS) without sacrificing diversity (comparable Precision/Recall). Notably, even without truncation, Re-MeanFlow matches MeanFlow trained from scratch on real data.
ImageNet $512^{2}$
Method FID $\downarrow$ IS $\uparrow$ Precision $\uparrow$ Recall $\uparrow$
Re-MeanFlow (no trunc.) 3.50 242.62 0.73 0.54
Re-MeanFlow (top $5\%$) 3.10 251.84 0.74 0.56